# JORA: JAX Tensor-Parallel LoRA Library for Retrieval Augmented Fine-Tuning

### 17 Mar 2024 v1

@@ Abstract
The scaling of Large Language Models (LLMs) for retrieval-based tasks, particularly in Retrieval Augmented Generation (RAG), faces significant memory constraints, especially when fine-tuning extensive prompt sequences. Current open-source libraries support full-model inference and fine-tuning across multiple GPUs but fall short of accommodating the efficient parameter distribution required for retrieved context. Addressing this gap, we introduce a novel framework for PEFT-compatible fine-tuning of Llama-2 models, leveraging distributed training. Our framework uniquely utilizes JAX’s just-in-time (JIT) compilation and tensor-sharding for efficient resource management, thereby enabling accelerated fine-tuning with reduced memory requirements. This advancement significantly improves the scalability and feasibility of fine-tuning LLMs for complex RAG applications, even on systems with limited GPU resources. Our experiments show more than 12x improvement in runtime compared to Hugging Face/DeepSpeed implementation with four GPUs while consuming less than half the VRAM per GPU. Our library will be open-sourced in due course.
@@
Anique Tahir Arizona State University 699 S. Mill Avenue Tempe, AZ research@anique.org                        Lu Cheng University of Illinois Chicago 851 S. Morgan St. Chicago, IL lucheng@uic.edu                        Huan Liu Arizona State University 699 S. Mill Avenue Tempe, AZ
@@ 5 An Example Usage Scenario
JORA is designed to aid in RAFT. In this section, we demonstrate a RAFT use case by fine-tuning it on a social media dataset. The purpose of RAG is to add additional context to a prompt by searching for knowledge and adding additional information. For RAFT, data can be created based on retrieved knowledge. The LLM learns to generate the retrieved answer based on the context since the key rationale is held back. A simple example is a database query, which corresponds to a process that may be taken to produce an output by evaluating the database. If the query is not provided but rather a natural language equivalent is provided, the LLM must learn the heuristics represented by the hidden query.
Since prompt tuning is insufficient for models to develop social understanding Gandhi et al. (2023), we use a fine-tuning process consisting of two phases to add knowledge to an LLM. Both phases of fine-tuning use PEFT. For our problem setting, rather than just predicting the following words, we aim to gain an understanding of the relation between different comments. For instance, a comment in a social media session may target the previous comment, the original post that spawned the session, or some comment in the middle of the discourse. To glean insight into the target of the comment in terms of its context, reasoning between the structure of the conversation is critical. Unfortunately, the LLM pre-training does not consider these relationships specifically, and there is no public data related to reasoning at the comment level in social media discourse. Thus, we rely on other general-purpose structured data as a surrogate to learn structure and reasoning. We use the WikiTableQuestions Pasupat and Liang (2015) dataset to infuse structural intelligence into the model. This dataset consists of various independent tables, questions based on one of the tables, and a corresponding answer. To answer these questions, using the data in the input table is vital. Some answers require aggregate reasoning.
@@
Table 2: The veracity of the directionality identification improves with the RAFT fine-tuning phases w.r.t. the baselines.
For the directionality analysis task (which post is targeted by another comment in the same session), we leveraged a corpus of 4chan threads Papasavva et al. (2020). This dataset consists of $\sim$3333 million threads and $\sim$100100100100 million posts. Because 4chan allows its users to tag whom they reply to, we use this data as the ground truth for directionality information. We examine whether our RAFT phases improve (i) the model’s ability to detect the post we are targeting for behavior comprehension and (ii) the model’s ability to distinguish who is being targeted by the poster. 4chan allows posters to mention more than one comment as the target of the reply. Here, we consider the model successful if one of the multiple comments is identified. Table 2 shows the result of our experiment. The RAFT model significantly improves performance over the pre-trained counterparts. This illustrates the application of RAFT to improve LLM performance in social media analysis. Social media conversation threads can provide important context but they can span large sequences. JORA helps in the training process here by splitting a discourse sequence’s computation tensors across multiple GPUs. This is not possible using HuggingFace/Deepspeed because Data-Parallelism in these frameworks distributes the workload between different data instances rather than dividing the computation for a single data instance among multiple processors.

### 19 Mar 2024 v2

@@ Abstract
The scaling of Large Language Models (LLMs) for retrieval-based tasks, particularly in Retrieval Augmented Generation (RAG), faces significant memory constraints, especially when fine-tuning extensive prompt sequences. Current open-source libraries support full-model inference and fine-tuning across multiple GPUs but fall short of accommodating the efficient parameter distribution required for retrieved context. Addressing this gap, we introduce a novel framework for PEFT-compatible fine-tuning of Llama-2 models, leveraging distributed training. Our framework uniquely utilizes JAX’s just-in-time (JIT) compilation and tensor-sharding for efficient resource management, thereby enabling accelerated fine-tuning with reduced memory requirements. This advancement significantly improves the scalability and feasibility of fine-tuning LLMs for complex RAG applications, even on systems with limited GPU resources. Our experiments show more than 12x improvement in runtime compared to Hugging Face/DeepSpeed implementation with four GPUs while consuming less than half the VRAM per GPU.
@@
Anique Tahir Arizona State University 699 S. Mill Avenue Tempe, AZ research@anique.org                        Lu Cheng University of Illinois Chicago 851 S. Morgan St. Chicago, IL lucheng@uic.edu                        Huan Liu Arizona State University 699 S. Mill Avenue Tempe, AZ huanliu@asu.edu
@@ 5 An Example Usage Scenario
JORA is designed to aid in RAFT. In this section, we demonstrate a RAFT use case by fine-tuning it on a social media dataset to help LLMs enable social-context understanding. The purpose of RAG is to add additional context to a prompt by searching for knowledge and adding additional information. For RAFT, data can be created based on retrieved knowledge. The LLM learns to generate the retrieved answer based on the context since the key rationale is held back. A simple example is a database query, which corresponds to a process that may be taken to produce an output by evaluating the database. If the query is not provided but rather a natural language equivalent is provided, the LLM must learn the heuristics represented by the hidden query.
Since prompt tuning is insufficient for models to develop social-context understanding Gandhi et al. (2023), we use a fine-tuning process consisting of two phases to add knowledge to an LLM. Both phases of fine-tuning use PEFT. For our problem setting, rather than just predicting the following words, we aim to gain an understanding of the relation across different comments in a social media session. For instance, a comment in a social media session may target the previous comment, the original post that spawned the session, or some comment in the middle of the discourse. To glean insight into the target of the comment in terms of its context, reasoning between the structure of the conversation is critical. Unfortunately, the LLM pre-training does not consider these relationships specifically, and there is no public data related to reasoning at the comment level in social media discourse. Thus, we rely on other general-purpose structured data as a surrogate to learn structure and reasoning. We use the WikiTableQuestions Pasupat and Liang (2015) dataset to infuse structural intelligence into the model. This dataset consists of various independent tables, questions based on one of the tables, and a corresponding answer. To answer these questions, using the data in the input table is vital. Some answers require aggregate reasoning.
@@
Table 2: The veracity of the directionality identification improves with the RAFT fine-tuning phases w.r.t. the baselines. Given the conversation as context, the values represent the accuracy of detecting the respective posts.
For the directionality analysis task (which post is targeted by another comment in the same session), we leveraged a corpus of 4chan threads Papasavva et al. (2020). This dataset consists of $\sim$3333 million threads and $\sim$100100100100 million posts. Because 4chan allows its users to tag whom they reply to, we use this data as the ground truth for directionality information. We examine whether our RAFT phases improve (i) the model’s ability to detect the post we are targeting for behavior comprehension and (ii) the model’s ability to distinguish who is being targeted by the poster. 4chan allows posters to mention more than one comment as the target of the reply. Here, we consider the model successful if one of the multiple comments is identified. Table 2 shows the result of our experiment. The RAFT model significantly improves performance over the pre-trained counterparts. This illustrates the application of RAFT to improve LLM performance in social media analysis. Social media conversation threads can provide important context but they can span large sequences. JORA helps in the training process here by splitting a discourse sequence’s computation tensors across multiple GPUs. This is not possible using HuggingFace/Deepspeed because Data-Parallelism in these frameworks distributes the workload between different data instances rather than dividing the computation for a single data instance among multiple accelerators.